# On the contraction properties of Sinkhorn semigroups

### v2

@@ 4.3 Weighted norm contractions
Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ holds as soon as condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied and the estimates (59) hold for some $n_{0}\geq 1$ and some parameters $\rho\in[0,1[$ and $\kappa>0$.
When ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied there also exist parameters $\kappa\geq 0$ and $\rho\in[0,1[$ that only depend on the parameters $(\epsilon,c)$ in (60) and on the function $\varepsilon(r)$ in (61) and such that, for any $n\geq n_{0}$,
$\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa},\psi_{\kappa}}({\cal S}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{\psi_{\kappa},\varphi_{\kappa}}({\cal S}_{2n+1})<\rho,\quad\text{where}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$
@@
Let us emphasize that Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is stronger than Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$. In particular, the Gibbs-loop transitions ${\cal S}^{\circ}_{n}$ can be stable even when the Sinkhorn transitions ${\cal S}_{n}$ are unstable. Intuitively, Gibbs-loop transitions go forward and backward towards their invariant measures (22) while the Sinkhorn transitions in (16) transport asymptotically a target marginal measure to the other. For instance, by Lemma 4.9 in [adm-24] (see also Section A) the Gibbs-loop transitions ${\cal S}^{\circ}_{n}$ of the Gaussian Sinkhorn algorithm are always stable. However, in the context of regularized Gaussian entropic problems on a time horizon $t$ (cf. Section 3.4 in [adm-24]) the Sinkhorn transitions ${\cal S}_{n}$ may be unstable when the time horizon is not sufficiently large (by Remarks 3.9 and 4.6 in [adm-24]).
4.4 A Lyapunov stability theorem
Regularity conditions
This section presents some sufficient conditions for ${\cal D}_{\varphi,\psi}({\cal S})$, in terms of the potential functions $(U,V,W)$. These simple conditions enable us to obtain geometric rates for the convergence of the Sinkhorn iterations for a variety of models, including polynomial growth models (Section 5.1) and heavy tailed models (Section 5.2), as well as boundary states models including Beta and Weibull distributions (Section 5.3 and Section 5.4). Illustrations in the context of Linear-Gaussian models and finite (non necessarily Gaussian) mixture marginals are also discussed respectively in Section 5.8 and Section 5.9.
@@ 4.4 A Lyapunov stability theorem
$W(x,y)\geq W_{\star}\quad\mbox{\rm for $(\lambda\otimes\nu)$-almost every $(x,y)\in(\mathbb{X}\times\mathbb{Y})$.}$ (63)
@@
$(e^{\delta U},e^{\delta V})\in({\cal B}_{\infty}(\mathbb{X})\times{\cal B}_{\infty}(\mathbb{Y})).$
@@
Condition 4.9 (${\cal H}_{\delta}(U,V)$).
There is some $\delta\in]0,1[$ such that the functions $({\cal U}_{\delta},{\cal V}_{\delta})$ defined in (27) are locally bounded with compact sub-level sets and we have
$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee~\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty.$ (64)
Note that (63) and (64) together imply the inequalities
@@
${\cal H}_{\delta}(U,V)\Longrightarrow\lambda_{(1-\delta)U}\left(1\right)=\lambda_{U}\left(\exp{(\delta U)}\right)<\infty\quad\mbox{\rm and}\quad\nu_{(1-\delta)V}(1)=\nu_{V}\left(\exp{(\delta V)}\right).$
In particular all absolute moments $\lambda_{U}(|U|^{p})\vee\nu_{V}(|V|^{p})<\infty$, with $p\geq 0$ are finite. Given the reference measure ${\cal P}_{0}=(\lambda_{U}\times{\cal Q})$ also have
${\cal H}_{\delta}(U,V)\Longrightarrow\mbox{\rm Ent}((\lambda_{U}\otimes\nu_{V})~|~{\cal P}_{0})=(\lambda_{U}\otimes\nu_{V})(W)-\nu_{V}(V)<\infty.$ (65)
@@
$\leq$ $W_{\star}+\lambda_{U}(e^{W^{V}-W_{\star}})=W_{\star}+e^{-W_{\star}}~\lambda_{(1-\delta)U}(e^{-{\cal U}_{\delta}}).$
@@
$\lambda_{(1-\delta)U}(e^{-{\cal U}_{\delta}})$ $\leq$ $e^{-r}~\lambda_{(1-\delta)U}(1)+\lambda_{(1-\delta)U}(e^{-{\cal U}_{\delta}}~1_{{\cal U}_{\delta}\leq r})$
@@
Condition 4.10 (${\cal H}^{\prime}_{\delta}(U,V)$).
There is some $\delta\in]0,1/2[$ such that ${\cal U}_{\delta}$ and ${\cal V}_{\delta}$ are locally bounded with compact sub-level sets and we have
$\lambda_{(1-2\delta)U}\left(1\right)\vee\nu_{(1-2\delta)V}\left(1\right)<\infty.$ (66)
We have
${\cal H}^{\prime}_{\delta}(U,V)\Longrightarrow{\cal H}_{\delta}(U,V).$ (67)
To verify this claim, note that the regularity condition on $({\cal U}_{\delta},{\cal V}_{\delta})$ ensures that
$(e^{{\cal U}_{\delta}},e^{{\cal V}_{\delta}})\in({\cal B}_{\infty}(\mathbb{X})\times{\cal B}_{\infty}(\mathbb{Y})).$
When ${\cal U}_{\delta}$ is locally bounded with compact sub-level sets $C_{{\cal U}}(r):=\{{\cal U}_{\delta}\leq r\}$ for any $0<\delta<1/2$ we have
$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)$ $=$ $\lambda_{(1-2\delta)U}\left(e^{-{\cal U}_{\delta}}\right)$
$\leq$ $\left(e^{-r}+e^{c_{\delta}(r)}\right)\lambda_{(1-2\delta)U}\left(1\right)\quad\mbox{\rm with}\quad c_{\delta}(r):=\sup_{x\in C_{{\cal U}}(r)}|{\cal U}_{\delta}(x)|.$
@@
This ends the proof of (67).
@@
Whenever $\|W\|<\infty$ we have $\|W_{U}\|\vee\|W^{V}\|<\infty$. In this case ${\cal H}_{\delta}(U,V)$ is satisfied as soon as there is some $\delta\in]0,1[$ such that
$\lambda_{(1-\delta)U}\left(1\right)\vee\nu_{(1-\delta)V}\left(1\right)<\infty.$ (69)
Remark 4.12.
Assume that $(U,V)$ and $(W_{U},W^{V})$ are continuous functions and we have $\delta U-W^{V}\geq{\cal U}_{\delta}^{-}$ and $\delta V-W_{U}\geq{\cal V}_{\delta}^{-}$, for some functions $({\cal U}_{\delta}^{-},{\cal V}_{\delta}^{-})$ with compact level sets. In this case, the functions $({\cal U}_{\delta},{\cal V}_{\delta})$ are locally bounded with compact level sets.
A Lyapunov stability theorem
@@
If Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is also satisfied with $n_{0}=1$ for the Lyapunov functions
$(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})\quad\mbox{with}\quad(U_{\star},V_{\star}):=(\inf_{\mathbb{X}}U,\inf_{\mathbb{Y}}V)\in(\mathbb{R}\times\mathbb{R}).$ (70)
@@
$\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|_{\psi}\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|_{\varphi}\leq c~\rho^{2n}\quad\text{for any $n\geq 1$.}$ (71)
@@
The above theorem combined with the estimates (62) ensures that
@@
with the parameters $(\epsilon,c)$ as in (60). The rescaling of the potential functions (70) is only made to have $\varphi\wedge\psi\geq 1$. Combining Theorem 4.13 with Theorem 4.8 we obtain the following corollary.
@@
Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. In this situation, there exists some parameters $n_{0}\geq 0$, $\rho\in[0,1[$ and $c_{1},c_{2}\geq 0$ such that for any $n\geq n_{0}$ we have
@@
The r.h.s. term in (72) can be estimate with (62). Also note that
@@
Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. In this situation, there exists some parameters $n_{0}\geq 0$, $\rho\in[0,1[$ and $c_{1},c_{2}\geq 0$ such that for any $n\geq n_{0}$ we have
@@
4.5 Wasserstein exponential decays
Assume that the Lyapunov functions $(\psi,\varphi)$ defined in (70) satisfy the estimates
$e^{\delta U(x_{1})}+e^{\delta U(x_{2})}\geq\varrho_{\varphi}(x_{1},x_{2})\quad\mbox{\rm and}\quad e^{\delta V(y_{1})}+e^{\delta V(y_{2})}\geq\varrho_{\psi}(y_{1},y_{2})$ (74)
for some semi-metric $\varrho_{\varphi},\varrho_{\psi}$. In this case, by (10) there exists $c_{u},c_{v}>0$ such that
@@ 4.5 Wasserstein exponential decays
with the Wasserstein semi-metrics $\mathbb{W}_{\varrho_{\varphi}},\mathbb{W}_{\varrho_{\psi}}$ defined in (8). This yields the following Wasserstein estimates.
@@
Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. In addition, the Lyapunov functions $(e^{\delta U},e^{\delta V})$ satisfy (74) for some semi-metric $(\varrho_{\varphi},\varrho_{\psi})$. In this situation, there exists some $n_{0}\geq 0$ and $c_{u},c_{v}\geq 0$ such that for any $n\geq n_{0}$ we have
@@
In the above display, $n_{0}\geq 1$ and $\rho\in[0,1[$ are as in (72) and $(\psi,\varphi)$ are the Lyapunov functions that depend on $\delta$ defined in (70).
@@
$V(y_{1})+V(y_{2})$ $\geq$ $c_{v}+\varrho_{V}(y_{1},y_{2}).$ (76)
In this case, by (35), for any $p\geq 1$ there exists some constant $c_{p}>0$ such that
@@
with the $p$-Wasserstein semi-metrics $\mathbb{W}_{p,\varrho_{U}},\mathbb{W}_{p,\varrho_{V}}$ defined in (8). This yields the following Wasserstein estimates.
@@
Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. In addition, the target potential functions $(U,V)$ satisfy (76) for some semi-metric $(\varrho_{U},\varrho_{V})$. In this situation, for any $p\geq 1$ there exists some $c_{p}\geq 0$ such that for any $n\geq n_{0}$ we have
@@
In the above display, $n_{0}\geq 1$ and $\rho\in[0,1[$ are as in (72) and $(\psi,\varphi)$ are the Lyapunov functions that depend on $\delta$ defined in (70).
@@
Using Proposition D.3 in Appendix D and following exactly the same argument as in the proof of Theorem 4.13, the estimate (71) remains valid when the weaker condition ${\cal D}_{\varphi,\psi}({\cal S})$ holds for some $n_{0}\geq 0$, provided that
@@
4.6 Entropy exponential decays
Let us discuss some fairly straightforward consequences of the estimate (71). First note that (71) ensures the convergence of the series
$\lim_{n\rightarrow\infty}U_{2n}=\mathbb{U}$ $:=$ $U_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}\quad\mbox{\rm in ${\cal B}_{\varphi}(\mathbb{X})$,}$
$\lim_{n\rightarrow\infty}V_{2n}=\mathbb{V}$ $:=$ $V_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p}}{d\nu_{V}}}\quad\mbox{\rm in ${\cal B}_{\psi}(\mathbb{Y})$},$ (78)
where $U_{0}=U$ and $V_{0}=0$. Applying Lebesgue’s dominated convergence theorem, the integral equations stated in (18) converge as $n\rightarrow\infty$ to a system of integral equations
$\mathbb{U}=U+\log{{\cal Q}(e^{-\mathbb{V}})}\quad\mbox{\rm and}\quad\mathbb{V}=V+\log{{\cal R}(e^{-\mathbb{U}})}.$
More precisely, using the estimates in (71) it is straightforward to prove the following corollary.

### v3

@@ 4.3. Weighted norm contractions
We have the following:
  • Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ holds as soon as condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ is satisfied and the estimates (4.12) hold for some $n_{0}\geq 1$ and some parameters $\rho\in[0,1[$ and $\kappa>0$.
  • When ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ is satisfied there exist parameters $\kappa\geq 0$ and $\rho\in[0,1[$ that only depend on the parameters $(\epsilon,c)$ in (4.13) and on the function $\varepsilon(r)$ in (4.14) and such that, for any $n\geq n_{0}$,
    $\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa},\psi_{\kappa}}({\mathcal{S}}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{\psi_{\kappa},\varphi_{\kappa}}({\mathcal{S}}_{2n+1})<\rho,\quad\text{where}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$
@@
Let us emphasize that Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ is stronger than Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$. In particular, the Gibbs-loop transitions ${\mathcal{S}}^{\circ}_{n}$ may be stable even in situations where the Sinkhorn transitions ${\mathcal{S}}_{n}$ are unstable. Intuitively, Gibbs-loop transitions alternate between forward and backward steps that steer the dynamics toward their invariant measures (2.8), whereas the Sinkhorn transitions in (2.2) asymptotically transport one target marginal measure to the other.
As an illustration, Lemma 4.9 in [adm-24] (see also Section A) shows that the Gibbs-loop transitions ${\mathcal{S}}^{\circ}_{n}$ of the Gaussian Sinkhorn algorithm are always stable. By contrast, in the setting of regularized Gaussian entropic problems over a finite time horizon $t$ (see Section 3.4 in [adm-24]), the Sinkhorn transitions ${\mathcal{S}}_{n}$ may become unstable when the time horizon is not sufficiently large; see Remarks 3.9 and 4.6 in [adm-24].
4.4. A Lyapunov stability theorem
4.4.1. Regularity conditions
This section presents some sufficient conditions for ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$, in terms of the potential functions $(U,V,W)$. These simple conditions enable us to obtain geometric rates for the convergence of the Sinkhorn iterations for a variety of models, including polynomial growth models (Section 5.1) and heavy tailed models (Section 5.2), as well as boundary states models including Beta and Weibull distributions (Section 5.3 and Section 5.4). Illustrations in the context of Linear-Gaussian models and finite (non necessarily Gaussian) mixture marginals are also discussed respectively in Section 5.8 and Section 5.9.
@@ 4.4.1. Regularity conditions
(4.16) $W(x,y)\geq W_{\star}\quad\mbox{\rm for $(\lambda\otimes\nu)$-almost every $(x,y)\in(\mathbb{X}\times\mathbb{Y})$.}$
@@
$(e^{\delta U},e^{\delta V})\in({\mathcal{B}}_{\infty}(\mathbb{X})\times{\mathcal{B}}_{\infty}(\mathbb{Y})).$
@@
Condition ${\mathcal{H}}_{\delta}(U,V)$: There is some $\delta\in]0,1[$ such that the functions $({\mathcal{U}}_{\delta},{\mathcal{V}}_{\delta})$ defined in (2.13) are locally bounded with compact sub-level sets and we have
(4.17) $\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee~\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty.$
Note that (4.16) and (4.17) together imply the inequalities
@@
${\mathcal{H}}_{\delta}(U,V)\Longrightarrow\lambda_{(1-\delta)U}\left(1\right)=\lambda_{U}\left(\exp{(\delta U)}\right)<\infty\quad\mbox{\rm and}\quad\nu_{(1-\delta)V}(1)=\nu_{V}\left(\exp{(\delta V)}\right).$
In particular all absolute moments $\lambda_{U}(|U|^{p})\vee\nu_{V}(|V|^{p})<\infty$, with $p\geq 0$ are finite. Given the reference measure ${\mathcal{P}}_{0}=(\lambda_{U}\times{\mathcal{Q}})$ also have
(4.18) ${\mathcal{H}}_{\delta}(U,V)\Longrightarrow\mbox{\rm Ent}((\lambda_{U}\otimes\nu_{V})~|~{\mathcal{P}}_{0})=(\lambda_{U}\otimes\nu_{V})(W)-\nu_{V}(V)<\infty.$
@@
$\leq$ $W_{\star}+\lambda_{U}(e^{W^{V}-W_{\star}})=W_{\star}+e^{-W_{\star}}~\lambda_{(1-\delta)U}(e^{-{\mathcal{U}}_{\delta}}).$
@@
$\lambda_{(1-\delta)U}(e^{-{\mathcal{U}}_{\delta}})$ $\leq$ $e^{-r}~\lambda_{(1-\delta)U}(1)+\lambda_{(1-\delta)U}(e^{-{\mathcal{U}}_{\delta}}~1_{{\mathcal{U}}_{\delta}\leq r})$
@@
Condition ${\mathcal{H}}^{\prime}_{\delta}(U,V)$: There is some $\delta\in]0,1/2[$ such that ${\mathcal{U}}_{\delta}$ and ${\mathcal{V}}_{\delta}$ are locally bounded with compact sub-level sets and we have
(4.19) $\lambda_{(1-2\delta)U}\left(1\right)\vee\nu_{(1-2\delta)V}\left(1\right)<\infty.$
Let us remark that
(4.20) ${\mathcal{H}}^{\prime}_{\delta}(U,V)\Longrightarrow{\mathcal{H}}_{\delta}(U,V).$
To verify this claim, note that the regularity condition on $({\mathcal{U}}_{\delta},{\mathcal{V}}_{\delta})$ ensures that
$(e^{{\mathcal{U}}_{\delta}},e^{{\mathcal{V}}_{\delta}})\in({\mathcal{B}}_{\infty}(\mathbb{X})\times{\mathcal{B}}_{\infty}(\mathbb{Y})).$
When ${\mathcal{U}}_{\delta}$ is locally bounded with compact sub-level sets $C_{{\mathcal{U}}}(r):=\{{\mathcal{U}}_{\delta}\leq r\}$ for any $0<\delta<1/2$ we have
$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)$ $=$ $\lambda_{(1-2\delta)U}\left(e^{-{\mathcal{U}}_{\delta}}\right)$
$\leq$ $\left(e^{-r}+e^{c_{\delta}(r)}\right)\lambda_{(1-2\delta)U}\left(1\right)\quad\mbox{\rm with}\quad c_{\delta}(r):=\sup_{x\in C_{{\mathcal{U}}}(r)}|{\mathcal{U}}_{\delta}(x)|.$
@@
This ends the proof of (4.20).
@@
Whenever $\|W\|<\infty$ we have $\|W_{U}\|\vee\|W^{V}\|<\infty$. In this case ${\mathcal{H}}_{\delta}(U,V)$ is satisfied as soon as there is some $\delta\in]0,1[$ such that
(4.22) $\lambda_{(1-\delta)U}\left(1\right)\vee\nu_{(1-\delta)V}\left(1\right)<\infty.$
4.4.2. A Lyapunov stability theorem
@@
If Condition ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ is also satisfied with $n_{0}=1$ for the Lyapunov functions
(4.23) $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})\quad\mbox{with}\quad(U_{\star},V_{\star}):=(\inf_{\mathbb{X}}U,\inf_{\mathbb{Y}}V)\in(\mathbb{R}\times\mathbb{R}).$
@@
(4.24) $\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|_{\psi}\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|_{\varphi}\leq c~\rho^{2n}\quad\text{for any $n\geq 1$.}$
@@
The above theorem combined with the estimates (4.15) ensures that
@@
with the parameters $(\epsilon,c)$ as in (4.13). The rescaling of the potential functions (4.23) is only made to have $\varphi\wedge\psi\geq 1$. Combining Theorem 4.8 with Theorem 4.6 we obtain the following corollary.
@@
Assume ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. In this situation, there exists some parameters $n_{0}\geq 0$, $\rho\in[0,1[$ and $c_{1},c_{2}\geq 0$ such that for any $n\geq n_{0}$ we have
@@
The r.h.s. term in (4.25) can be estimate with (4.15). Also note that
@@
Assume ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. In this situation, there exists some parameters $n_{0}\geq 0$, $\rho\in[0,1[$ and $c_{1},c_{2}\geq 0$ such that for any $n\geq n_{0}$ we have
@@
Assume that $(U,V)$ and $(W_{U},W^{V})$ are continuous functions and we have
$\delta U-W^{V}\geq{\mathcal{U}}_{\delta}^{-}\quad\mbox{\rm and}\quad\delta V-W_{U}\geq{\mathcal{V}}_{\delta}^{-},$
for some functions $({\mathcal{U}}_{\delta}^{-},{\mathcal{V}}_{\delta}^{-})$ with compact level sets. In this case, the functions $({\mathcal{U}}_{\delta},{\mathcal{V}}_{\delta})$ are locally bounded with compact level sets. As underlined in [ajay-dp-25], using (C.7) and Proposition D.9, the continuity condition in the above assertion can be relaxed. More precisely, Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$, Theorem 4.8 as well as Corollaries  4.9 and 4.10 hold when $({\mathcal{U}}_{\delta}^{-},{\mathcal{V}}_{\delta}^{-})$ are locally bounded with compact sub-level sets. Assume we have
$W(x,y)\leq a_{W}(x)+b_{W}(y)$
for some locally bounded functions $(a_{W},b_{W})$ such that $\lambda_{U}(a_{W})\vee\nu_{V}(b_{W})<\infty$. Note that
${\mathcal{U}}^{-}_{\delta}(x):=\delta U(x)-a_{W}(x)-\nu_{V}(b_{V})\leq{\mathcal{U}}_{\delta}(x)\leq\delta U(x)-W_{\star}.$
with $W_{\star}$ as in (4.16). This shows that $({\mathcal{U}}_{\delta},{\mathcal{U}}^{-}_{\delta})$ and similarly $({\mathcal{V}}_{\delta},{\mathcal{V}}^{-}_{\delta})$ with ${\mathcal{V}}^{-}_{\delta}:=\delta V-b_{W}$ are locally bounded. Moreover $({\mathcal{U}}^{-}_{\delta},{\mathcal{V}}^{-}_{\delta})$ have compact compact sub-level sets if and only if $(\delta U-a_{W},\delta V-b_{W})$ have compact compact sub-level sets.
4.5. Wasserstein exponential decays
Assume that the Lyapunov functions $(\psi,\varphi)$ defined in (4.23) satisfy the estimates
(4.27) $e^{\delta U(x_{1})}+e^{\delta U(x_{2})}\geq\varrho_{\varphi}(x_{1},x_{2})\quad\mbox{\rm and}\quad e^{\delta V(y_{1})}+e^{\delta V(y_{2})}\geq\varrho_{\psi}(y_{1},y_{2})$
for some semi-metric $\varrho_{\varphi},\varrho_{\psi}$. In this case, by (1.10) there exists $c_{u},c_{v}>0$ such that
@@ 4.5. Wasserstein exponential decays
with the Wasserstein semi-metrics $\mathbb{W}_{\varrho_{\varphi}},\mathbb{W}_{\varrho_{\psi}}$ defined in (1.8). This yields the following Wasserstein estimates.
@@
Assume ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. In addition, the Lyapunov functions $(e^{\delta U},e^{\delta V})$ satisfy (4.27) for some semi-metric $(\varrho_{\varphi},\varrho_{\psi})$. In this situation, there exists some $n_{0}\geq 0$ and $c_{u},c_{v}\geq 0$ such that for any $n\geq n_{0}$ we have
@@
In the above display, $n_{0}\geq 1$ and $\rho\in[0,1[$ are as in (4.25) and $(\psi,\varphi)$ are the Lyapunov functions that depend on $\delta$ defined in (4.23).
@@
(4.29) $V(y_{1})+V(y_{2})$ $\geq$ $c_{v}+\varrho_{V}(y_{1},y_{2}).$
In this case, by (3.5), for any $p\geq 1$ there exists some constant $c_{p}>0$ such that
@@
with the $p$-Wasserstein semi-metrics $\mathbb{W}_{p,\varrho_{U}},\mathbb{W}_{p,\varrho_{V}}$ defined in (1.8). This yields the following Wasserstein estimates.
@@
Assume ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. In addition, the target potential functions $(U,V)$ satisfy (4.29) for some semi-metric $(\varrho_{U},\varrho_{V})$. In this situation, for any $p\geq 1$ there exists some $c_{p}\geq 0$ such that for any $n\geq n_{0}$ we have
@@
In the above display, $n_{0}\geq 1$ and $\rho\in[0,1[$ are as in (4.25) and $(\psi,\varphi)$ are the Lyapunov functions that depend on $\delta$ defined in (4.23).
@@
Using Proposition D.3 in Appendix D and following exactly the same argument as in the proof of Theorem 4.8, the estimate (4.24) remains valid when the weaker condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ holds for some $n_{0}\geq 0$, provided that
@@
4.6. Entropy exponential decays
Let us discuss some fairly straightforward consequences of the estimate (4.24). First note that (4.24) ensures the convergence of the series
$\lim_{n\rightarrow\infty}U_{2n}=\mathbb{U}$ $:=$ $U_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}\quad\mbox{\rm in ${\mathcal{B}}_{\varphi}(\mathbb{X})$,}$
(4.31) $\lim_{n\rightarrow\infty}V_{2n}=\mathbb{V}$ $:=$ $V_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p}}{d\nu_{V}}}\quad\mbox{\rm in ${\mathcal{B}}_{\psi}(\mathbb{Y})$},$
where $U_{0}=U$ and $V_{0}=0$. Applying Lebesgue’s dominated convergence theorem, the integral equations stated in (2.4) converge as $n\rightarrow\infty$ to a system of integral equations
$\mathbb{U}=U+\log{{\mathcal{Q}}(e^{-\mathbb{V}})}\quad\mbox{\rm and}\quad\mathbb{V}=V+\log{{\mathcal{R}}(e^{-\mathbb{U}})}.$
More precisely, using the estimates in (4.24) it is straightforward to prove the following corollary.